OpenStudy (anonymous):

int ((e^x)^2/(e^x)^2+3e^x +2)dx

OpenStudy (anonymous):

\[\int\limits \frac{ (^2e^x) }{ (e^x)^2 +3e^x +2}\]

hartnn (hartnn):

split the numerator e^x^2 as e^x times e^x and put u = e^x du= ... ?